# arXiv:2603.01211v1[cs.AI] 01 Mar 2026

### v1

@@ 5.3 Critical Ethical Considerations
Defining the Boundary of Measurement: Finally, it is important to be clear about the scope of our claim: we are measuring the cultural intelligence of a model, not determining when it can be considered truly intelligent. Determining “intelligence” requires a level of additional grounding that goes beyond our current scope of measurement. This includes establishing which specific cultural context a model should adhere to, when that adherence is appropriate, and, critically, how to ethically aggregate the diverse results (including those tied to sensitive or contradictory norms) into a holistic, actionable verdict.
@@
Trust and the Illusion of Safety: Another major ethical risk lies in how the measurement results are interpreted by the end-user. There is a tangible risk that users will uncritically trust high cultural intelligence scores, assuming the results imply a system is perfectly safe, comprehensive, and appropriate for their context. This can lead to unwarranted reliance on the system’s judgments about other cultures. We must clearly communicate that these measures represent a limited assessment of capabilities, enabling to track progress of capabilities over time, and should not be mistaken for a comprehensive guarantee of safety or cultural competence in real-world, high-stakes scenarios.
@@
Navigating Contested and Contradictory Norms: Furthermore, an evaluation framework that relies on cultural norms must acknowledge that these practices are often contested. Within communities, norms are constantly being challenged, recreated through resistance, and redefined through social and political movements. An evaluation that rigidly assesses adherence to existing, dominant norms may be undesirable if it risks reinforcing an oppressive status quo. Compounding this, cultural norms and values may sometimes be in direct contradiction with one another or with the core values upheld by the model developers or platforms. Our knowledge-based approach offers a key ethical lever here: by making the underlying norms explicit, it allows model builders to selectively adhere to desirable norms and strategically override harmful ones, an ethical intervention that purely perception-based evaluations cannot easily afford.
@@
The sociopolitical nature of cultural data: While incorporating community-based data engagement is crucial in our proposed framework, such data and its use are inherently political [46, 34]. The challenge of mitigating these risks is further compounded by the unique complexities associated with social identities, historical narratives, and power dynamics [28]. They can be socially situated and context-dependent [46], evolving over time [7], and subject to diverse interpretations within and across cultural groups [9]. Furthermore, cultural data may encompass sensitive information or traditional knowledge that requires careful handling to respect cultural values, privacy, and intellectual property rights [85]. For our measurements to be responsible, it must move beyond simply collecting data to establishing ethical, inclusive partnerships with communities. These partnerships ensure that the data used to define and measure AI capabilities are socially situated, context-dependent, and respectful of the sensitive information and traditional knowledge that requires careful handling to protect cultural values and intellectual property rights. This is especially important in the context of indigenous cultures, where centuries of exploitation and injustice have led to a deep-seated distrust communities and corporations [19], while recent efforts have been actively working to prioritize Indigenous knowledge and community leadership [24, 31].

### v2

@@ 5.3 Critical Ethical Considerations
The scope of our effort is in measuring the cultural intelligence of a model, not determining when it can be considered truly intelligent. Determining “intelligence” requires a level of additional grounding that goes beyond our current scope of measurement. This includes establishing which specific cultural context a model should adhere to, when that adherence is appropriate, and, critically, how to ethically aggregate the diverse results (including those tied to sensitive or contradictory norms) into a holistic, actionable verdict. Another major ethical risk lies in how the measurement results are interpreted by the users. There is a tangible risk that users will uncritically trust high cultural intelligence scores, assuming the results imply a system is perfectly safe, comprehensive, and appropriate for their context. This can lead to unwarranted reliance on the system’s judgments about other cultures. These measures represent a limited assessment of capabilities, enabling to track progress of capabilities over time, and should not be mistaken for a comprehensive guarantee of safety or cultural competence in real-world, high-stakes scenarios.
@@
Furthermore, within communities, various cultural facets, such as norms are constantly being challenged, recreated through resistance, and redefined through social and political movements. An evaluation that rigidly assesses adherence to existing, dominant norms may be undesirable if it risks reinforcing an oppressive status quo. Our knowledge-based approach offers a key ethical lever here: by making the underlying norms explicit, it allows model builders to selectively adhere to desirable norms and strategically override harmful ones, an ethical intervention that purely perception-based evaluations cannot easily afford.
@@
While incorporating community-based data engagement is crucial in our proposed framework, such data and its use are inherently political [89]. The challenge of mitigating these risks is further compounded by the unique complexities associated with social identities, historical narratives, and power dynamics [90]. They can be socially situated and context-dependent [89], evolving over time [91], and subject to diverse interpretations within and across cultural groups [92]. Furthermore, cultural data may encompass sensitive information or traditional knowledge that requires careful handling to respect cultural values, privacy, and intellectual property rights [93]. Hence, it is important to move beyond simply collecting data to establishing ethical, inclusive partnerships with communities.